# Look-ahead Search on Top of Policy Networks in Imperfect Information Games

### v2

@@ Acknowledgments
This research is supported by Czech Science Foundation (grant no. 18-27483Y) and Grant Agency of the Czech Technical University in Prague, grant No. SGS22/168/OHK3/3T/13. Computational resources were supplied by the project ”e-Infrastruktura CZ” (e-INFRA CZ LM2018140) supported by the Ministry of Education, Youth and Sports of the Czech Republic and also the OP VVV funded project CZ.02.1.01/0.0/0.0/16_019/0000765 “Research Center for Informatics” which are both gratefully acknowledged.
@@ Appendix A Proofs
This proof is heavily based on the proof of the original V-trace [Espeholt et al., 2018], with the notable difference being that we discuss only the proof for finite horizon with discount factor $\gamma=1$. We can reformulate the V-trace operator as
@@ C.1 Counterfactual values in gadget games
(a) Leduc hold’em
(b) Goofspiel 5 descending
@@ C.2 Predefined transformations
(a) Goofspiel 5 descending
(b) Goofspiel 5 randomized
@@ Appendix D Additional experimental details
We used the same setting for all the hyperparameters for all the experiments. The setting is shown in Table 2. For the majority of hyperparameters, we adopted the settings established by the original authors in their work on RNaD [Perolat et al., 2022]. The differences are setting importance sampling clipping $\overline{\rho}$ to $\infty$ to ensure that the estimate of V-trace is unbiased. Furthermore, we changed the rate of updating the regularization policy, learning rate, batch size, and size of hidden layers to compensate for training on smaller games than the original authors did with Stratego. As a sampling policy $\mu$, we have used the actor policy, effectively avoiding importance sampling clipping by working on-policy.
@@
Table 2: Parameters used during evaluation of SeRNaD

### v3

@@ Acknowledgments
This research is supported by Czech Science Foundation (GA22-26655S) and Grant Agency of the CTU in Prague (SGS22/168/OHK3/3T/13), and the Canadian Institute for Advanced Research. Computational resources were supplied by (e-INFRA CZ LM2018140) supported by the Ministry of Education, Youth and Sports of the Czech Republic and also (CZ.02.1.01/0.0/0.0/16 019/0000765).
@@ Appendix A Proofs
This proof of Theorem 1 is based on the proof of the original V-trace Espeholt et al. [2018], with the notable difference that we discuss only the case with finite horizon and discount factor $\gamma=1$. We can reformulate the V-trace operator as
@@ C.1 Counterfactual values in gadget games
((a)) Leduc hold’em
((b)) Goofspiel 5 descending
@@ C.2 Predefined transformations
((a)) Goofspiel 5 descending
((b)) Goofspiel 5 randomized
@@ Appendix D Additional experimental details
We used the same setting for all the hyperparameters for all the experiments. The setting is shown in Table 2. For the majority of hyperparameters, we adopted the settings established by the original authors in their work on RNaD Perolat et al. [2022]. The differences are setting importance sampling clipping $\overline{\rho}$ to $\infty$ to ensure that the estimate of V-trace is unbiased. Furthermore, we changed the rate of updating the regularization policy, learning rate, batch size, and size of hidden layers to compensate for training on smaller games than the original authors did with Stratego. As a sampling policy $\mu$, we have used the actor policy, effectively avoiding importance sampling clipping by working on-policy.
@@
Table 2: Parameters used during evaluation of SePoT
@@ Appendix E Pseudocodes
Appendix F SePoT components diagram
SePoT consists of multiple components that interact together. In Figure 6 we show these components and their connection. The process of the SePoT can be summarized as follows: During training, the trajectories are sampled from the environment. These are then used to train the policy network with the underlying policy gradient algorithm. We generate the transformations for the same trajectories based on the policy network and train the transformation network with them. All three of policy and transformations from networks and trajectories are then used to train the estimate the expected value for each transformation, which is used for training the critic. This process is repeated for a specified amount of training iterations. During gameplay, when a player is about to make a decision, it either decides to sample action based on the policy network or refine the strategy with search. When refining the strategy, the algorithm first reconstructs the public state, which is all the histories that share the same public knowledge. Then, it uses a policy network to compute the reach of each infoset. After that, the depth-limited subgame is constructed, which requires computing counterfactual values at the root and giving the opponent a choice between different strategies after the depth limit. For both of these, we use the trained critic. This depth-limited subgame is then solved with some search algorithm; in our case, it was CFR+.
Figure 6: Main components of SePoT